OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

so comparing with y = mx+c slope of y = 4x-2 is 4. Now what would be the slope of a line parallel to this one?

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

Ok there is a concept. "Two lines, in co-ordinate geometry are said to be parallel to each other if they have EQUAL SLOPES." So now u can tell slope of parallel line?

OpenStudy (anonymous):

if they have equal slopes it would be 4

OpenStudy (anonymous):

yes. so now slope of new line M = 4, and it passes through (-3, 6) Let (-3, 6) be (x1, y1) respectively. Then equation of line is given by = y - y1 = m(x-x1).